# Improving Robustness to Model Inversion Attacks via Sparse Coding Architectures

### v1

@@ Abstract
Recent model inversion attack algorithms permit adversaries to reconstruct a neural network’s private training data just by repeatedly querying the network and inspecting its outputs. In this work, we develop a novel network architecture that leverages sparse-coding layers to obtain superior robustness to this class of attacks. Three decades of computer science research has studied sparse coding in the context of image denoising, object recognition, and adversarial misclassification settings, but to the best of our knowledge, its connection to state-of-the-art privacy vulnerabilities remains unstudied. However, sparse coding architectures suggest an advantageous means to defend against model inversion attacks because they allow us to control the amount of irrelevant private information encoded in a network’s intermediate representations in a manner that can be computed efficiently during training and that is known to have little effect on classification accuracy. Specifically, compared to networks trained with a variety of state-of-the-art defenses, our sparse-coding architectures maintain comparable or higher classification accuracy while degrading state-of-the-art training data reconstructions by factors of $1.1$ to $18.3$ across a variety of reconstruction quality metrics (PSNR, SSIM, FID). This performance advantage holds across $5$ datasets ranging from CelebA faces to medical images and CIFAR-10, and across various state-of-the-art SGD-based and GAN-based inversion attacks, including Plug-&-Play attacks. We provide a cluster-ready PyTorch codebase to promote research and standardize defense evaluations.
@@ 1 Introduction
The popularization of machine learning has been accompanied by the widespread use of neural networks that were trained on private, sensitive, and proprietary datasets. This has given rise to a new generation of privacy attacks that seek to infer private information about the training dataset simply by inspecting the representation of the training data that remains encoded in the model’s parameters [20, 21, 37, 93, 49, 80, 88, 31, 89, 64, 68, 12, 44].
@@
Such attacks are feasible because each hidden layer of a standard network architecture captures a detailed representation of the training data. It is well-known that standard dense layers exhibit a tendency to memorize their inputs [23, 10, 59], so even a minimal leak of a network’s class distribution output or a leak of its intermediate outputs from a single layer is often sufficient to train an inverse mapping for data reconstruction. More concretely, state-of-the-art inversion attacks work by submitting externally obtained images to the model, observing leaked outputs, then using this data to train a new ‘inverted’ neural network that reconstructs (predicts) an input image given a leaked output. This can be accomplished either directly via SGD, or by optimizing a GAN, and we consider both approaches here. Such attacks on standard network architectures can reconstruct private training images that are clearly recognizable by humans familiar with the training data [30, 87, 26, 83, 4, 35, 68, 22].
@@
Very little is known about how a network’s architecture contributes to its robustness (or vulnerability). This is surprising since throughout three decades of research in other domains such as image denoising [6, 19, 13, 54, 9, 60, 42, 3], object recognition [55, 65, 39, 24], and adversarial misclassification [69, 56, 40, 70], researchers seeking to control their model’s representations of the data have heavily studied sparse coding-based architectures that prune unnecessary details and preserve only the information that is essential to the model objective. Specifically, sparse coding seeks to approximately represent an image (or layer) with only a small set of basis vectors selected from an overcomplete dictionary [19, 54, 9]. While it is well-known that computing a sparse representation using a standard objective function is NP-hard in general [53, 16, 33], we now benefit from fast approximation algorithms that efficiently compute high-quality sparse representations [43, 60, 39, 42, 33, 52, 8, 14]. Sparse coding architectures leverage this technique by inserting a sparse network layer after a dense layer, such that the sparse layer reduces the dense layer’s outputs to a sparse representation.
@@
Main contribution. We begin by showing that an off-the-shelf sparse coding-based architecture offers performance advantages compared to state-of-the-art data augmentation, regularization, and noise based defenses in terms of robustness to model inversion attacks. We then refine this idea to achieve superior performance. Our main result is a novel sparse-coding based architecture, SCA, that is robust to state-of-the-art model inversion attacks.
@@ 2 Threat models
3. Split network threat model (Federated Learning). We also consider the split network threat model described by [71]. There has been much recent interest in Federated Learning architectures that split the network across multiple agents [41, 48, 7], particularly for privacy-fraught domains such as medicine where legal requirements limit data sharing [73, 36]. These architectures are known to be susceptible to model inversion attacks,[71], and defenses are urgently needed. This threat model also allows us to capture a different view of a ‘worst-case’ threat model: Model inversion attacks are known to be more effective when the attacker has access to outputs from earlier layers that may exhibit a more direct representation of the input images [26]. To capture this ‘worst-case’, we consider the setting where the attacker has access to raw intermediate outputs from the first linear network layer. As before, we also assume the attacker has access to ideal training data examples drawn from the actual training datasets.
@@ 3 SCA architecture
Among different techniques to perform sparse coding, we leverage the commonly used Locally Competitive Algorithm (LCA) [60]. LCA implements a recurrent network of leaky integrate-and-fire neurons that incorporates the general principles of thresholding and feature-similarity-based competition between neurons to solve Equation 1. While Rozell introduced LCA in the non-convolutional setting, it can be readily adapted to the convolutional setting (see Section 0.A.2 for details). Specifically, each LCA neuron has an internal membrane potential $\mathcal{P}$ which evolves per the following differential equation:
@@
where $\tau$ is a time constant, $\Psi(t)=\mathcal{X}*\Omega$ is the neuron’s bottom-up drive from the input computed by taking the convolution, ***, between the input, $\mathcal{X}$, and the dictionary, $\Omega$, and $-\mathcal{P}(t)$ is the leak term [70, 40]. Lateral competition between neurons is performed via the term $-\mathcal{R}_{x}(t)*\mathcal{G}$, where $\mathcal{G}=\Omega*\Omega-I$ is the similarity between each feature and the other $\mathcal{F}$ features ($-I$ prevents self interactions). $\mathcal{R}_{x}$ is computed by applying soft threshold activation $T_{\lambda}(x)=\textnormal{relu}(x-\lambda)$ to the neuron’s membrane potential, which produces nonnegative, sparse representations. Overall, this means that in LCA neurons will compete to determine which ones best represent the input, and thus will have non-zero activations in $\mathcal{R}_{x}$, the output of the SCL that is passed to the next layer.
@@
The SCA architecture is defined by the use of multiple pairs of sparse coding and dense (batch norm) layers after the input image, which can then be followed by other (linear, convolutional) layers. Fig. 2 illustrates this design principle. The key intuition is that the first sparse layer jettisons unnecessary private information in the input image. Then, by alternating sparse-dense pairs of layers, we ensure that unnecessary information is also jettisoned from downstream layers. In this manner, downstream layers also do not convey unnecessary private information to the adversary, and they also do not e.g. learn to reconstruct private information jettisoned by the first sparse layer. In short, previous defenses work by trying to mislead attackers by pushing model features in a wrong direction, either randomly via noise or strategically via adversarial examples, or by disincentivizing memorization during training. In contrast, SCA directly removes the unnecessary private information.
Training SCA is identical to training a standard network with one exception: after each backprop updates non-sparse layers, we perform a fast update on the sparse layers, except for the first sparse layer that sparse-codes the image input. [^1] We can optionally also allow backpropagation to update this very first sparse layer after the input image. We do this in our experiments. Alternatively, in some learning scenarios it may be advantageous to instead precompute the sparse representation of each image and delete the original images before training, as the first sparse layers remain fixed when we optionally exclude them from backpropagation.
@@
While we focus on the neuron lateral competition approach to sparse coding as it is practically convenient and well-represented in recent work [70], we note that for large-scale machine learning applications, we now have practical parallel algorithms that learn the sparse coding dictionary near-optimally w.p. in parallel time (adaptivity) that is logarithmic in the size of the data [33, 8, 14]. Fast single-iteration heuristics are also available (see e.g. [84]). Thus, even for large-scale applications, computing sparse representations while training SCA adds little computational overhead compared to sophisticated optimization-based techniques necessary for recent defenses [22]. In practice, even our basic sparse coding research implementations (see Section 4 and Appendix 0.A.3 below) are slightly faster than optimized Torch implementations of the best-performing recent defense [58].
@@ 4 Experiments
Our goal in this section is to show that SCA performs well compared to state-of-the-art defenses as well as practical defenses used in leading industry models in terms of both classification accuracy and various reconstruction quality metrics. To evaluate its performance comprehensively, we test all combinations of $5$ diverse datasets, $3$ threat models, $9$ defense baselines, plus multiple runs-per-experiment and various sparsity parameters $\lambda$. Appendices 0.A.1-0.A.6 give details.
@@
For each dataset, we conduct three sets of experiments corresponding to our three threat models. First, we test SCA and baselines’ defenses against a state-of-the-art Plug-&-Play attack [68] that leverages StyleGAN3 to obtain high-quality reconstructions. Second, we compare SCA networks to a variety of baselines in terms of their robustness to a state-of-the-art end-to-end network attack that leverages leaked raw high-dimensional outputs from the networks’ last hidden layer, as well as held-out training data drawn from the true training dataset. This allows us to assess SCA’s defenses in a realistic setting with a well-informed adversary. Our third set of experiments tests performance in a split network setting of [71] where the attacker has access to leaked raw outputs from the first linear network layer. Robustness in this setting is desirable because model inversion attacks are known to be more effective on earlier hidden layers [26], and also because an algorithm that is robust to such attacks would be an effective defense under novel security paradigms such as Federated Learning, which is known to be vulnerable to inversion [71]. Appendix 0.A.5 gives details.
@@
  • Titcombe et al. [71]. We train a state-of-the-art Laplace $\mathcal{L}(\mu$$=$$0,b$$=$0.5)0.5)0.5 ) noise defense as in [71]. We also try more noise—see Appendix 11 and 12
@@
In all experiments, we consider the simplest case of SCA architecture that contains SCA’s alternating sparse-and-dense layer pairs followed by only linear layers. We note that adding downstream convolutional layers or more sophisticated downstream architectures is certainly possible, though we avoid this here in order to compare the essence of the SCA approach to the baselines. Appendix 0.A.4 describes SCA architecture details. In the split network setting, we are careful to use slightly shallower SCA architectures with fewer linear layers to match the split network experiments of [71].
@@
    $FID^{2}({X_{in}},{X^{*}_{in}})=||\mu_{X_{in}}-\mu_{X^{*}_{in}}||^{2}+Tr(\mathrm{Cov}_{X_{in}}+\mathrm{Cov}_{X^{*}_{in}}-2*\sqrt{\mathrm{Cov}_{X_{in}}\cdot\mathrm{Cov}_{X^{*}_{in}}})$
@@
For all experiments, we consider the standard train test split of $70$% and $30$%. After training each defense model, we run attacks to reconstruct the entire training set and compare reconstruction performance. We run all the experiments on a standard industry production cluster with $4$ nodes and DELL Tesla V100 GPUs with $40$ cores. We provide a cluster-ready (anonymized) PyTorch codebase that implements attacks, SCA, other defenses, and replication codes for all experiments at: https://github.com/SayantonDibbo/SCA.
@@ 4.1 Experimental results overview
Across the $3$ attack settings and $5$ datasets, SCA maintains comparable or higher classification accuracy while degrading state-of-the-art training data reconstructions $1.1$ to $11.7$ times more than the $9$ baselines in terms of PSNR and FID, and $1.1$ to $720$ times more in terms of SSIM. This performance gap exceeds the scale of improvements made by of recent algorithms, and SCA’s defense advantage is also more stable than alternatives across multiple runs. This is because unlike for baselines, even an idealized ‘perfect attacker’ can only hope to recover a sparsified, un-detailed training image from SCA. We show results here for the $2$ most privacy-sensitive datasets of medical images and CelebA faces under all $3$ threat models, and defer the $9$ other {dataset, attack} combinations to Appendix 0.C.
Figure 3: Qualitative comparisons among actual and reconstructed images (Plug-&-Play Attack [68]) under SCA and baselines on hi-res CelebA dataset.

### v2

@@ Abstract
Recent model inversion attack algorithms permit adversaries to reconstruct a neural network’s private and potentially sensitive training data by repeatedly querying the network. In this work, we develop a novel network architecture that leverages sparse-coding layers to obtain superior robustness to this class of attacks. Three decades of computer science research has studied sparse coding in the context of image denoising, object recognition, and adversarial misclassification settings, but to the best of our knowledge, its connection to state-of-the-art privacy vulnerabilities remains unstudied. In this work, we hypothesize that sparse coding architectures suggest an advantageous means to defend against model inversion attacks because they allow us to control the amount of irrelevant private information encoded by a network in a manner that is known to have little effect on classification accuracy. Specifically, compared to networks trained with a variety of state-of-the-art defenses, our sparse-coding architectures maintain comparable or higher classification accuracy while degrading state-of-the-art training data reconstructions by factors of $1.1$ to $18.3$ across a variety of reconstruction quality metrics (PSNR, SSIM, FID). This performance advantage holds across $5$ datasets ranging from CelebA faces to medical images and CIFAR-10, and across various state-of-the-art SGD-based and GAN-based inversion attacks, including Plug-&-Play attacks. We provide a cluster-ready PyTorch codebase to promote research and standardize defense evaluations.
@@ 1 Introduction
The popularization of machine learning has been accompanied by the widespread use of neural networks that were trained on private, sensitive, and proprietary datasets. This has given rise to a new generation of privacy attacks that seek to infer private information about the training dataset simply by inspecting the representation of the training data that remains encoded in the model’s parameters [23, 24, 40, 98, 52, 84, 93, 34, 94, 68, 72, 13, 19, 47, 18].
@@
Such attacks are feasible because each hidden layer of a standard network architecture captures a detailed representation of the training data. It is well-known that standard dense layers exhibit a tendency to memorize their inputs [26, 11, 63], so even a minimal leak of a network’s class distribution output or a leak of its intermediate outputs from a single layer is often sufficient to train an inverse mapping for data reconstruction. More concretely, state-of-the-art inversion attacks work by submitting externally obtained images to the model, observing leaked outputs, then using this data to train a new ‘inverted’ neural network that reconstructs (predicts) an input image given a leaked output. This can be accomplished either directly via SGD, or by optimizing a GAN, and we consider both approaches here. Such attacks on standard network architectures can reconstruct private training images that are clearly recognizable by humans familiar with the training data [33, 92, 29, 87, 4, 38, 72, 25, 21].
@@
Very little is known about how a network’s architecture contributes to its robustness (or vulnerability). This is surprising since throughout three decades of research in other domains such as image denoising [6, 22, 14, 58, 10, 64, 45, 3], object recognition [59, 69, 42, 27], and adversarial misclassification [73, 60, 43, 74, 20], researchers seeking to control their model’s representations of the data have heavily studied sparse coding-based architectures that prune unnecessary details and preserve only the information that is essential to the model objective. Specifically, sparse coding seeks to approximately represent an image (or layer) with only a small set of basis vectors selected from an overcomplete dictionary [22, 58, 10]. While it is well-known that computing a sparse representation using a standard objective function is NP-hard in general [56, 17, 36], we now benefit from fast approximation algorithms that efficiently compute high-quality sparse representations [46, 64, 42, 45, 36, 55, 8, 15]. Sparse coding architectures leverage this technique by inserting a sparse network layer after a dense layer, such that the sparse layer reduces the dense layer’s outputs to a sparse representation.
@@
Main contribution. We begin by showing that an off-the-shelf sparse coding preprocessing step offers performance advantages compared to state-of-the-art data augmentation, regularization, and noise based defenses in terms of robustness to model inversion attacks. We then refine this idea into a network architecture that achieves superior performance. Our main result is a novel sparse-coding architecture, SCA, that is robust to state-of-the-art model inversion attacks.
@@ 2 Threat models
3. Split network threat model (Federated Learning). We also consider the split network threat model described by [75]. There has been much recent interest in Federated Learning architectures that split the network across multiple agents [44, 51, 7, 21, 29], particularly for privacy-fraught domains such as medicine where legal requirements limit data sharing [77, 39]. These architectures are known to be susceptible to model inversion attacks,[75, 29, 21], and defenses are urgently needed. This threat model also allows us to capture a different view of a ‘worst-case’ threat model: Model inversion attacks are known to be more effective when the attacker has access to outputs from earlier layers that may exhibit a more direct representation of the input images [29]. To capture this ‘worst-case’, we consider the setting where the attacker has access to raw intermediate outputs from the first linear network layer. As before, we also assume the attacker has access to ideal training data examples drawn from the actual training datasets. Appendix B provides additional details about the model inversion threat models.
@@ 3 SCA architecture
Among different techniques to perform sparse coding, we leverage the commonly used Locally Competitive Algorithm (LCA) [64]. LCA implements a recurrent network of leaky integrate-and-fire neurons that incorporates the general principles of thresholding and feature-similarity-based competition between neurons to solve Equation 1. While Rozell introduced LCA in the non-convolutional setting, it can be readily adapted to the convolutional setting (see Appendix A.2 for details). Specifically, each LCA neuron has an internal membrane potential $\mathcal{P}$ which evolves per the following differential equation:
@@
where $\tau$ is a time constant, $\Psi(t)=\mathcal{X}*\Omega$ is the neuron’s bottom-up drive from the input computed by taking the convolution, *, between the input, $\mathcal{X}$, and the dictionary, $\Omega$, and $-\mathcal{P}(t)$ is the leak term [74, 43]. Lateral competition between neurons is performed via the term $-\mathcal{R}_{x}(t)*\mathcal{G}$, where $\mathcal{G}=\Omega*\Omega-I$ is the similarity between each feature and the other $\mathcal{F}$ features ($-I$ prevents self interactions). $\mathcal{R}_{x}$ is computed by applying soft threshold activation $T_{\lambda}(x)=\textnormal{relu}(x-\lambda)$ to the neuron’s membrane potential, which produces nonnegative, sparse representations. Overall, this means that in LCA neurons will compete to determine which ones best represent the input, and thus will have non-zero activations in $\mathcal{R}_{x}$, the output of the SCL that is passed to the next layer.
@@
The SCA architecture is defined by the use of multiple pairs of sparse coding and dense (batch norm) layers after the input image, which can then be followed by other (linear, convolutional) layers. Fig. 2 illustrates this design principle. The key intuition is that the first sparse layer jettisons unnecessary private information in the input image. Then, by alternating sparse-dense pairs of layers, we ensure that unnecessary information is also jettisoned from downstream layers. In this manner, downstream layers also do not convey unnecessary private information to the adversary, and they also do not e.g. learn to reconstruct private information jettisoned by the first sparse layer. In short, previous defenses work by trying to mislead attackers by pushing model features in a wrong direction, either randomly via noise or strategically via adversarial examples, or by disincentivizing memorization during training. In contrast, SCA directly removes the unnecessary private information. Training SCA is identical to training a standard network with one exception: after each backprop updates non-sparse layers, we perform a fast update on the sparse layers, except for the first sparse layer that sparse-codes the image input. [^1] We can optionally also perform a backpropagation on sparse layers after updating them each iteration via sparse coding. We do this in our experiments.
@@
While we focus on the neuron lateral competition approach to sparse coding as it is practically convenient and well-represented in recent work [74], we note that for large-scale machine learning applications, we now have practical parallel algorithms that learn the sparse coding dictionary near-optimally w.p. in parallel time (adaptivity) that is logarithmic in the size of the data [36, 8, 15]. Fast single-iteration heuristics are also available (see e.g. [88]). Thus, even for large-scale applications, computing sparse representations while training SCA adds little computational overhead compared to sophisticated optimization-based techniques necessary for recent defenses [25]. In practice, even our basic sparse coding research implementations (see Section 4 and Appendix A.3) are slightly faster than optimized Torch implementations of the best-performing recent defense [62].
@@ 4 Experiments
xOur goal in this section is to show that SCA performs well compared to state-of-the-art defenses as well as practical defenses used in leading industry models in terms of both classification accuracy and various reconstruction quality metrics. To evaluate its performance comprehensively, we test all combinations of $5$ diverse datasets, $3$ threat models, $9$ defense baselines, plus multiple runs-per-experiment and various sparsity parameters $\lambda$. See Appendix A.1-A.6.
@@
For each dataset, we conduct three sets of experiments corresponding to our three threat models. First, we test SCA and baselines’ defenses against a state-of-the-art Plug-&-Play attack [72] that leverages StyleGAN3 to obtain high-quality reconstructions. Second, we compare SCA networks to a variety of baselines in terms of their robustness to a state-of-the-art end-to-end network attack that leverages leaked raw high-dimensional outputs from the networks’ last hidden layer, as well as held-out training data drawn from the true training dataset. This allows us to assess SCA’s defenses in a realistic setting with a well-informed adversary. Our third set of experiments tests performance in a split network setting of [75] where the attacker has access to leaked raw outputs from the first linear network layer. Robustness in this setting is desirable because model inversion attacks are known to be more effective on earlier hidden layers [29], and also because an algorithm that is robust to such attacks would be an effective defense under novel security paradigms such as Federated Learning, which is vulnerable to inversion [75] (see Appendix A.5).
@@
  • Titcombe et al. [75]. We train a state-of-the-art Laplace $\mathcal{L}(\mu$$=$$0,b$$=$0.5)0.5)0.5 ) noise defense as in [75]. We also try more noise—see Appendix D.
@@
In all experiments, we consider the simplest case of SCA architecture that contains SCA’s alternating sparse-and-dense layer pairs followed by only linear layers. We note that adding downstream convolutional layers or more sophisticated downstream architectures is certainly possible, though we avoid this here in order to compare the essence of the SCA approach to the baselines. Appendix A.4 describes SCA details. In the split network setting, we are careful to use slightly shallower SCA architectures with fewer linear layers to match the split network experiments of [75].
@@
    $FID^{2}({X_{in}},{X^{*}_{in}})=||\mu_{X_{in}}-\mu_{X^{*}_{in}}||^{2}+Tr(\mathrm{Co}_{X_{in}}+\mathrm{Co}_{X^{*}_{in}}-2*\sqrt{\mathrm{Co}_{X_{in}}\cdot\mathrm{Co}_{X^{*}_{in}}})$
@@
For all experiments, we consider the standard train test split of $70$% and $30$%. After training each defense model, we run attacks to reconstruct the entire training set and compare reconstruction performance. We run all the experiments on a standard industry production cluster with $4$ nodes and DELL Tesla V100 GPUs with $40$ cores. We provide a cluster-ready PyTorch codebase on our project page at: https://sayantondibbo.github.io/SCA.
@@ 4.1 Experimental results overview
Defense. Across the $3$ attack settings and $5$ datasets, SCA maintains comparable or higher classification accuracy while degrading state-of-the-art training data reconstructions $1.1$ to $11.7$ times more than the $9$ baselines in terms of PSNR & FID, and $1.1$ to $720$ times more in terms of SSIM. This performance gap exceeds the scale of improvements made by recent algorithms. SCA’s defense is also more stable than baselines across multiple runs. This is because unlike for baselines, even an idealized ‘perfect attacker’ can only hope to recover a sparsified, un-detailed training image from SCA. We show results here for the $2$ most privacy-sensitive datasets of medical images and CelebA in $3$ threat models, and defer the $9$ other {dataset, attack} combinations to Appendix C.
Accuracy. Typically, obtaining greater defense means trading away accuracy. However, in $6$ of the $15$ experiments (MNIST + FashionMNIST)$\times$(Plug & Play + end-to-end + split networks), SCA outperforms no-defense and all baselines’ accuracy. SCA also outperforms all baselines’ accuracy on CelebA in Plug & Play, and a sparse approach is within 0.003 of the best accuracy on an 8th experiment. No other single SOTA baseline wins on accuracy this consistently. We emphasize that unlike baselines that do accuracy hyperparameter tuning, we obtain this result absent any such tuning. SCA drops accuracy on MedMNIST (which is the most imbalanced & has fewest training examples). However, tuning of SCA (kernel size 5 $\rightarrow$ 7) improves SCA’s accuracy on MedMNIST in Table 3 from 94.6% to 97%—See Appendix G and Table 12.
Figure 3: Experiments set 1: Qualitative comparisons among actual and reconstructed images (Plug-&-Play Attack [72]) under SCA & baselines on hi-res CelebA dataset.
@@
Basic Sparse-Standard outperforms SOTA baselines. Our basic Sparse-Standard baseline outperforms the best baselines’ PSNR on both CelebA and Medical MNIST, and also outperforms baselines’ SSIM on the latter. SCA0.5 then outperforms Sparse-Standard on all metrics, obtaining SSIM and FID that are better by factors of $1.2$ to $1.6$ and $1.1$ to $1.2$, respectively, and slightly better PSNR. Thus, while Sparse-Standard offers an inferior defense vs. SCA, it still offers a fast practical defense for less privacy-critical domains.
@@
CelebA No-Defense $11.35$ $0.718$ $256.4$ $0.779$
Gaussian-Noise $10.39$ $0.604$ $264.8$ $0.644$
GAN $10.15$ $0.613$ $289.7$ $0.635$
Titcombe et al.[75] $10.18$ $0.636$ $304.1$ $0.654$
Gong et al.[25]++ $10.02$ $0.556$ $381.5$ $0.672$
Gong et al.[25] $10.11$ $0.595$ $350.5$ $0.614$
Peng et al.[62] $9.90$ $0.514$ $402.8$ $0.728$
Hayes et al.[28] $9.92$ $0.556$ $383.7$ $0.621$
Wang et al.[82] $9.85$ $0.527$ $402.8$ $0.742$
Sparse-Standard $9.84$ $0.539$ $374.7$ $0.728$
SCA0.1 $\mathbf{9.79}$ $\mathbf{0.451}$ $\mathbf{391.9}$ $\mathbf{0.726}$
SCA0.25 $\mathbf{9.45}$ $\mathbf{0.442}$ $\mathbf{411.0}$ $\mathbf{0.739}$
SCA0.5 $\mathbf{9.40}$ $\mathbf{0.440}$ $\mathbf{412.6}$ $\mathbf{0.723}$
Medical No-Defense $22.04$ $0.396$ $196.1$ $0.998$
MNIST Gaussian-Noise $21.83$ $0.382$ $209.4$ $0.862$
GAN $21.77$ $0.427$ $219.0$ $0.998$
Gong et al.[25]++ $21.50$ $0.359$ $273.1$ $0.894$
Titcombe et al.[75] $21.68$ $0.360$ $286.3$ $0.899$
Gong et al.[25] $21.75$ $0.477$ $249.1$ $0.770$
Peng et al.[62] $21.82$ $0.381$ $268.3$ $0.927$
Hayes et al.[28] $21.72$ $0.337$ $259.7$ $0.823$
Wang et al.[82] $21.71$ $0.322$ $211.7$ $0.937$
Sparse-Standard $20.97$ $0.086$ $239.3$ $0.907$
SCA0.1 $\mathbf{21.19}$ $\mathbf{0.057}$ $\mathbf{253.5}$ $\mathbf{0.888}$
SCA0.25 $\mathbf{21.17}$ $\mathbf{0.075}$ $\mathbf{280.1}$ $\mathbf{0.882}$
SCA0.5 $\mathbf{20.06}$ $\mathbf{0.055}$ $\mathbf{288.8}$ $\mathbf{0.881}$
Figure 4: Experiments set 2: Qualitative comparisons among actual & reconstructed images (end-to-end setting) under SCA & baselines on the Medical MNIST dataset.